OpenStudy (chrisplusian):

Need some help understanding properties of functions defined by power series

OpenStudy (anonymous):

Do you have a specific question?

OpenStudy (chrisplusian):

The general form they give for integeration of a power series is:\[\int\limits f(x)dx=C+\sum_{n=0}^{\infty}a _{n}\frac{ (x-c)^{n+1} }{ n+1 }=c+a_{0}(x-c)+a _{1}(x-c)^{2}+....\] how do I apply something like that to a generic nth term?

OpenStudy (chrisplusian):

ex f(x)=\[\sum_{n=1}^{\infty}\frac{ x ^{n} }{ n }\]

OpenStudy (chrisplusian):

how would I find the anti-derivative of this function?

OpenStudy (chrisplusian):

would it be just generically......\[\int\limits f(x)dx=\sum_{n=1}^{\infty}\frac{ x ^{n+1} }{ n(n+1) }\]

OpenStudy (anonymous):

@amistre64 could you help

OpenStudy (amistre64):

yes

OpenStudy (chrisplusian):

the part I don't get is where in their general form they have the a sub nth term of the sequence times the anti-derivative.

OpenStudy (amistre64):

since the only x parts in the power series come for the exponentiated "x" the rest is just constants

OpenStudy (chrisplusian):

sorry it shoudl have said that I don't understand why they have the a sub nth term of the SERIES times the anti derivative

OpenStudy (amistre64):

\[f(x)=\sum_k c_nx^n\] \[\int f(x)dx=\int\sum_k c_nx^n~dx\] \[F(x)+c=\sum_k \frac{c_n}{n+1}x^{n+1}\]

OpenStudy (amistre64):

all a series is is a polynomial of so many terms; integrating a sum is equal to integrating each term

OpenStudy (chrisplusian):

So then do I need to multiply the anti-derivative times the original a sub nth term?

OpenStudy (amistre64):

of course there are certain convergence issues to address, but overall ..

OpenStudy (amistre64):

a_n is a constant ...

OpenStudy (chrisplusian):

ok so then it would just be using properties of a series where you can multiply a constant to a series after the series is evaluated?

OpenStudy (chrisplusian):

Then the original radius of convergence still applies but I would have to evaluate the end points to determine if they are in the interval of convergence?

OpenStudy (chrisplusian):

I would have to evaluate them in the antiderivative series to see if they converge or diverge right?

OpenStudy (amistre64):

\[f(x)=\sum c_nx^n=c_0+c_1x+c_2x^2+...+c_nx^n+...\] \[\int f(x)=\int(c_0+c_1x+c_2x^2+...+c_nx^n+...)\] \[\int f(x)=\int c_0+\int c_1x+\int c_2x^2+\int ...+\int c_nx^n+\int ...\]

OpenStudy (amistre64):

im not read up on how convergences effect it, but for the most part, you can just assume it until yelled at later :)

OpenStudy (chrisplusian):

lol thanks

OpenStudy (amistre64):

yw